OpenStudy (anonymous):

Hi everyone! Here is a nasty rectangular to polar conversion, 4x^2-5y^2-36y-36=0...can anyone help solve for r? Thanks!

OpenStudy (anonymous):

I have tried sooo many combinations and still can't get it

OpenStudy (primeralph):

singing calc2 blues? lol

OpenStudy (anonymous):

:o)

OpenStudy (primeralph):

.......just plug in rsint and rcost......

OpenStudy (anonymous):

i'm not creating parametric equations...just polar

OpenStudy (primeralph):

by t I mean theta............

OpenStudy (anonymous):

is your t meant to be a theta?

OpenStudy (anonymous):

\[x= r \cos \theta\] \[y= r \sin \theta\] \[4x^2-5y^2-36y-36=0\] \[4(r \cos \theta)^2-5(r \sin \theta)^2-36r \sin \theta-36=0\] \[4r^2 \cos^2 \theta-5r^2 \sin^2 \theta-36r \sin \theta-36=0\] \[(4 \cos^2 \theta -5 \sin^2 \theta) r^2 -(36 \sin \theta) r - 36 =0\] from here i would assume is just a messy quadratic equation

OpenStudy (anonymous):

i've to that point before...i just can't solve for r

OpenStudy (primeralph):

same story...you can solve for r by using the quad formula

OpenStudy (anonymous):

i can?

OpenStudy (anonymous):

i mentioned messy right \[r= \frac{-b \pm \sqrt{b^2- 4ac}}{2a}\]

OpenStudy (primeralph):

yes

OpenStudy (primeralph):

..............but it's gonna look ugly along the line........

OpenStudy (anonymous):

unless you can find an nice transformation that i cant remember off the top of my head good luck

OpenStudy (primeralph):

the final answer is quite neat..........

OpenStudy (anonymous):

there are like 4 solutions it could be, and none of them are ugly...

OpenStudy (primeralph):

I said...ugly along the line

OpenStudy (anonymous):

like -4/1+sin theta is an example...i just am having serious trouble getting towards that form

OpenStudy (anonymous):

i will try the quadratic thingy...omg...this is gonna take awhile...can you periodically check in on me to see if i get it?

OpenStudy (primeralph):

okay, I can see you're smart, so Imma help you out. \[\frac{ -6 }{ 3\sin(\theta) \pm 2}\]

OpenStudy (anonymous):

woa...one of my choices as an answer is 6/2-3sin theta...that looks a little different than your plus or minus thingy...thingy is a technical term fyi :o)

OpenStudy (anonymous):

ok...i'm gonna try and work this AGAIN...it might take awhile though...:o)

OpenStudy (primeralph):

duhhh, that means there are two solutions, one with + and the other with -. lol

OpenStudy (anonymous):

can i ask you a question?

OpenStudy (primeralph):

if you bring the - from the 6 down, it swaps the position of the 2 and 3sint

OpenStudy (primeralph):

ask...........

OpenStudy (anonymous):

what tipped you off that you should be using the quadratic formula? Is it kinda just one of those things where you sit back and recognize the bigger picture?

OpenStudy (primeralph):

yeah........kinda. Once I saw an r and an r^2.........that meant quad.

OpenStudy (anonymous):

neat...okay...i'm gonna try and work this out...brb...well not right back lol

OpenStudy (primeralph):

yo, I already gave you the answer....and you seen it already in the options. I'm leaving and won't be able to confirm whatever you do if it takes longer than 2 minutes.

OpenStudy (anonymous):

no problem...thanks so much for your help Prime!

OpenStudy (primeralph):

Thanks. Tag me whenever......

OpenStudy (anonymous):

ok :o)